OpenStudy (anonymous):

Find the maximun y-value on the graph of y=f(x). f(x)=-x^2+2x+2

OpenStudy (tkhunny):

You must select words more carefully. Words mean things. You have defined y = f(x). Therefore, it makes no sense to sak "What is y?" You just defined it. If you copied this accurately, you should complain to the author. Anyway, how do you propose that we might find a solution? Perhaps a 1st Derivative? Maybe a 2nd Derivative? What say you?

OpenStudy (anonymous):

change -x^2 +2x+2 into completed square form

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

no how do I do that

OpenStudy (tkhunny):

So, algebra it is, then. You MUST know how to complete the square. Utilize this principle: \((a+b)^{2} =a^{2} +2ab+b^{2}\) How do you apply that to your right-hand expression?

OpenStudy (anonymous):

notsu re

OpenStudy (anonymous):

-x^2+2x+2 -(x^2-2x)+2 -{(x)^2 - 2*x*1 + 1^2 - 1^2 } +2 -(x-1)^2 +1+2 -(x-1)^2+3

OpenStudy (anonymous):

So, f(x)=-(x-1)^2+3 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Now, (x-1)^2>=0 u know it?

OpenStudy (anonymous):

is that y

OpenStudy (anonymous):

y=f(x)=-(x-1)^2+3

OpenStudy (anonymous):

so y=(x-1)^2+3

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

U missed the -ve sign

OpenStudy (anonymous):

where

OpenStudy (anonymous):

y=-(x-1)^2+3

OpenStudy (anonymous):

oh gotsa

OpenStudy (anonymous):

So, (x-1)^2>=0 -(x-1)^2<=0 -(x-1)^2+3<=3 y<=3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So, the max value of y is 3 got it?

OpenStudy (anonymous):

yes